Weather: Highs in the upper 40’s. A mix of calm to light winds. Cloudy to partly cloudy conditions throughout the day. It was nice not having to wear sunglasses for a change. Current NWS forecast:

Some excitement this morning! 

While John was troubleshooting the RASS which wasn't working when we arrived (see John’s ISS blog post for details), I visited the SP site which was just down the road to check the soil sensors and measure an HRXL reference height (228.6 cm).


Soils sensors remain buried but the soil looks porous. I saw no glaring issues.

Doing a walk through I noticed that one of the ratched cables was off the rebar so I re-hooked and tightened it (had a devil of a time figuring out how to loosen the strap on the ratched). I also tightened one of the other cables which was flopping in the breeze. 


Back at ISS1 Isabel texted that the DSM was off the net (great timing while we were at ISS1 !!). With Isabel’s guidance, I power cycled the DSM which re-started NIDAS but not the connection. Isabel suggested checking the ethernet cables - so I unplugged/replugged the ethernet cables in the DSM box, while John checked the ethernet attachment in the trailer. It looks like the ethernet in the trailer was strained and not well connected. John loosened and re-attached the ethernet more securely. Everything so far is green and working at ISS1. 

John's troubleshooting got the RASS working and will closely monitor it (again refer to John’s ISS blog for details). The Profiler/RASS has been offline from ~ 1020a - 1150a. 

Launched the 415p radiosonde. 

  • No labels

1 Comment

  1. On the Ethernet cable, the cable is an outdoor cable and is very rigid.  The cable runs up through the hole in the counter and into an Ethernet switch mounted on the wall.   As a result of the cable's rigidity, it was being pulled laterally out of its socket on the switch.  I pulled in some extra cable and zipped tied it to one of the eye bolts on the counter to align and provide strain relief for the cable's connector into the switch.